# Testing Quantumness via Photon Statistics for Time-Bin based Quantum Random Number Generators

### v1

@@ Abstract
where $\mu$ is the expected number of photons per unit time. A generic block diagram of the photon arrival time-based QRNG is given in Figure 1.
@@
Figure 2: Time diagram for the arrivals of the photons. There are three different methods to generate random numbers from photon arrival times. (I) Photons are detected in a continuous time reference at time $\{t_{0},t_{1},\dots t_{2(i-1)},t_{2i-1},t_{2i},\dots\}$. The $i$-th bit of the random number will be 00 if $\Delta t_{2i-1}<\Delta t_{2i}$ and $1$ if $\Delta t_{2i-1}>\Delta t_{2i}$. (II) Time is divided into bins with equal lengths. If a photon has been detected in the $i$-th bin then $i$ will be the corresponding random number. A new cycle of bins starts after a fixed sleep time ($t_{s}$) of the SPD. (III) An external time reference of cycle $T$ has been used. Each cycle is divided into $N$ equal time-bins $\tau_{1},\tau_{2},\dots,\tau_{N}$. For each cycle, if the first photon has been detected in $i$-th bin then $i$ will be the corresponding random number. SPD will be in sleep mode for the remaining time of the cycle.
There are different methods to produce random numbers from the registered times. Time diagrams for these methods are shown in Figure 2. In Figure 2(I), if the registered times of photon arrivals are $\{t_{0},t_{1},\dots,t_{2i-2},t_{2i-1},t_{2i},\dots\}$, the non-overlapping consecutive time intervals, i.e., $\Delta t_{2i-1}=t_{2i-1}-t_{2i-2},\Delta t_{2i}=t_{2i}-t_{2i-1}$, are used to generate random numbers as follows. If $\Delta t_{2i-1}<\Delta t_{2i}$, the $i$-th bit of the random number is taken as 00. If $\Delta t_{2i-1}>\Delta t_{2i}$, the $i$-th bit is taken as $1$. But in this process the rate of the generation of random number is slow. The length of the random bit-string is half of the photon count. Comparatively fast processes have been developed in the literature Michael A. Wayne and Kwiat (2009); Wayne and Kwiat (2010); Nie et al. (2014). In Figure 2(II), the time is divided into time-bins, each having equal length. If a photon is detected in the $i$-th time-bin, $i$ is taken as the random number. A new cycle of the bins starts after a fixed sleep time of the detector. Since the photon counting process is a Poisson process, the probability, that the waiting time ($t_{w}$) between the arrivals of two consecutive photons is greater than $t$, is given by $P(t_{w}>t)=P(\text{no photon in }(0,t])=P(0,t)=e^{-\mu t}$. Therefore, the random number generated with this procedure follows exponential distribution Michael A. Wayne and Kwiat (2009); Wayne and Kwiat (2010). To overcome this drawback, a new scheme is proposed in Nie et al. (2014). In this scheme, as depicted in Figure 2(III), an external time reference is used. This external time is divided into cycles of length $T$, which is smaller than the dead time of the SPD. The time $T$ is further divided into $N$ equal subintervals, called the time-bins $\tau_{1},\tau_{2},\dots,\tau_{N}$. In each cycle, if the first photon is detected in the $i$-th time-bin, $i$ would be the corresponding random number. SPD would be in sleep mode for the remaining time of the cycle.
@@
Photon statistics for testing quantumness with ideal devices. We first discuss the actual photon statistics in the practical scenario for external reference time-based QRNG (Figure 2(III)). Here, a random number is generated based on the detection time of the first photon in each time cycle $T$. The probability of detecting the first photon in $i$-th bin, when $n$ photons are present within $T$ time, is given by
@@
Photon statistics for testing quantumness with erroneous devices. Till now, we have considered all of the used optical devices as ideal. But in practice, any error in the source and in the attenuators affects photon power and changes the value of the mean photon number per unit time. Again, if the detector is not $100\%$ efficient, The first photon would be detected in the $i$-th time-bin, if at least one photon has been detected in $i$-th bin, and either (A) no photon has been reached at SPD before the $i$-th bin, or (B) $k$ number of photons have reached before the $i$-th bin, but none of them has been detected. Therefore, the probability of detecting the first photon in the $i$-th bin, when $n$ photons are present within $T$ time, is given by
@@
Since in the probability distribution of Equation (8), $\frac{T}{N}$ is the length of the time-bins, the corresponding probability mass function is also given by Equation (8) with $T=Nt_{i}$.
@@
$0.05$ $13$-$14/15$
$0.10$ $11$-$13/15$
$0.15$ $10$-$11/15$
$0.20$ $9$-$10/15$
$0.25$ $7$-$9/15$
$0.30$ $7$-$8/15$
$0.35$ $6$-$8/15$
$0.40$ $6$-$7/15$
Table 1: NIST test result for the data simulated with the distribution given by Equation (8). Three files of size $1$ GB have been used for the test for every $\mu Td$. The data for $\mu Td=0.05$ has passed all but the universal statistical test.
Therefore, these two types of QRNGs perform similarly in the practical scenarios. However, if we can ensure that the source is an actual single-photon source, then the perfect random number can be generated using an external time reference with probability distribution $P_{d}(i)=\frac{d}{N}$ (see Equation (3)). Since the distributions given by Equation (8) are evaluated from the quantum properties of photons, this can be used to test the quantumness in a QRNG. For this testing, one may use the $\chi^{2}$ goodness-of-fit test Dodge (2008) on the data generated by QRNGs as follows. Consider null and alternative hypotheses as
H0: The data follow the distribution (8).
[…]
The test statistic is defined as $\chi^{2}=\sum_{i=1}^{N}\frac{(O_{i}-E_{i})^{2}}{E_{i}}$, where $O_{i}$ denotes the frequency of $i$ in the data, $E_{i}=nf(i)$, $n$ is the size of generated data. Then the null hypothesis will be rejected if $\chi^{2}>\chi^{2}_{N-1,\alpha}$, where $\chi^{2}_{N-1,\alpha}$ is the value given in the chi-square table for significance level $\alpha$. Note that, the typical values of the parameters $\mu,T,d,N$ in $f(i)$ are known to the QRNG designer.
$P_{d}(1;1st)-P_{d}(N;1st)<\epsilon$
$\implies e^{-\frac{\mu Td}{N}}+e^{-(N-1)\frac{\mu Td}{N}}-e^{-\mu Td}>1-\epsilon.$ (11)
We have generated random numbers simulating the probability distribution given by Equation (8) for different $\mu Td$. Three files of size $1$ GB have been generated for every $\mu Td$. Then we perform the NIST tests over the generated numbers. $100$ bit streams of length $100,000$ have been considered for the test. There are a total of $15$ tests in the NIST testing suite. A perfectly random number should pass all of them. In Table 1, We tabulate the number of tests our simulated data passed. The data for $\mu Td=0.05$ has passed all but the universal statistical test. This table indicates that lower $\mu Td$ values give rise to random numbers that require less post-processing.
Conclusion. In this letter, we discuss photon statistics for time-bin-based quantum random number generation for practical scenarios. We consider two photon arrival time-based QRNG models, one using external time reference and another without such external reference. We show that, although for ideal devices these two models follow different distributions (one of which is perfectly random), these two QRNGs follow a similar distribution when the photon source is not an actual single photon source. We also suggest a test for quantumness using the $\chi^{2}$ goodness-of-fit test. If the data passes the test, one can say that the data has been generated using the quantum properties. We derive a relation when the photon statistics becomes $\epsilon$-random. We found that by making the product of the expected number of photons ($\mu T$) within an external reference cycle ($T$) and the detector efficiency ($d$) small enough, the randomness of the generated raw random number can be increased. Therefore it would require less post-processing, and it would retain maximum quantumness. In practical cases, the dark count of a detector also contributes to the generation of random numbers. In our discussion, we did not consider the dark count. This work can be extended to see the effect of dark counts on the photon statistics. One may note that our statistics are not meant for black-box testing for commercially available QRNGs. Rather, the physicists and engineers may use these statistics in the QRNG design phase before post-processing, for evaluation, fine-tuning and performance optimization.
Acknowledgement. S. Das would like to acknowledge the support received by the Quantum Delta NL KAT-2 project.

### v2

@@ IV.2 QRNG based on Coherent Pulse
A generic block diagram of the photon arrival time-based QRNG using photons from coherent sources is given in Figure 3.
SourceAttenuatorSPDFPGARandom Number
Figure 3: A basic block diagram of the photon arrival time-based QRNG using weak coherent pulses. Coherent pulses, coming from the source, are sent through attenuators to reduce their power so that the probability of having multiple photons within a predefined time interval is negligible Argillander et al. (2023). A SPD detects these photons and the detection time (photon arrival time) has been registered. The FPGA provides random numbers based on the registered time stamps.
@@
Figure 4: Time diagram for the arrivals of the photons. There are three different methods to generate random numbers from photon arrival times. (I) Photons are detected in a continuous time reference at time $\{t_{0},t_{1},\dots t_{2(i-1)},t_{2i-1},t_{2i},\dots\}$. The $i$-th bit of the random number will be 00 if $\Delta t_{2i-1}<\Delta t_{2i}$ and $1$ if $\Delta t_{2i-1}>\Delta t_{2i}$. (II) Time is divided into bins of equal length. If a photon has been detected in the $i$-th bin then $i-1$ will be the corresponding random number. A new cycle of bins starts after a fixed sleep time ($t_{s}$) of the SPD. (III) An external time reference of cycle $T$ has been used. Each cycle is divided into $N$ equal time-bins $\tau_{1},\tau_{2},\dots,\tau_{N}$. For each cycle, if the first photon has been detected in $i$-th bin then $i-1$ will be the corresponding random number. SPD will be in sleep mode for the remaining time of the cycle.
There are different methods to produce random numbers from the registered times. Time diagrams for these methods are shown in Figure 4. In Figure 4(I), if the registered times of photon arrivals are $\{t_{0},t_{1},\dots,t_{2i-2},t_{2i-1},t_{2i},\dots\}$, the non-overlapping consecutive time intervals, i.e., $\Delta t_{2i-1}=t_{2i-1}-t_{2i-2},\Delta t_{2i}=t_{2i}-t_{2i-1}$, are used to generate random numbers as follows. If $\Delta t_{2i-1}<\Delta t_{2i}$, the $i$-th bit of the random number is taken as 00. If $\Delta t_{2i-1}>\Delta t_{2i}$, the $i$-th bit is taken as $1$. But in this process the random number generation rate is slow. The length of the random bit-string is half of the photon count. Comparatively fast processes have been developed in the literature Michael A. Wayne and Kwiat (2009); Wayne and Kwiat (2010); Nie et al. (2014). In Figure 4(II), the time is divided into time-bins, each having equal length. If a photon is detected in the $i$-th time-bin, $i-1$ is taken as the random number. A new cycle of the bins starts after the detector has a fixed sleep time. Since the photon counting process is a Poisson process, the probability, that the waiting time ($t_{w}$) between the arrivals of two consecutive photons is greater than $t$, is given by $P(t_{w}>t)=P(\text{no photon in }(0,t])=P(0,t)=e^{-\mu t}$. Therefore, the random number generated with this procedure follows exponential distribution Michael A. Wayne and Kwiat (2009); Wayne and Kwiat (2010). To overcome this drawback, a new scheme is proposed in Nie et al. (2014). In this scheme, as depicted in Figure 4(III), an external time reference is used. This external time is divided into cycles of length $T$, which is smaller than the dead time of the SPD. The time $T$ is further divided into $N$ equal subintervals, called the time-bins $\tau_{1},\tau_{2},\dots,\tau_{N}$. In each cycle, if the first photon is detected in the $i$-th time-bin, $i-1$ would be the corresponding random number. SPD would be in sleep mode for the remaining time of the cycle.
@@
IV.3 Time-bin Statistics for Poissonian Photon with Ideal Devices
Currently available photon arrival time-based QRNGs use coherent sources, which generate photons that follow Poissonian statistics. Here we discuss the actual photon statistics in the practical scenario for external reference time-based QRNG (Figure 4(III)) to see the randomness of the generated random number. Here, a random number is generated based on the detection time of the first photon in each time cycle $T$. The probability of detecting the first photon in $i$-th bin, when $n$ photons are present within $T$ time, is given by
@@ IV.3 Time-bin Statistics for Poissonian Photon with Ideal Devices
IV.4 Time-bin Statistics for Poissonian Photon with Erroneous Devices
Till now, we have considered all of the used optical devices as ideal. But in practice, any error in the source and in the attenuators affects photon power and changes the value of the mean photon number per unit time. Again, if the detector is not $100\%$ efficient, The first photon would be detected in the $i$-th time-bin, if at least one photon has been detected in $i$-th bin, and either (A) no photon has been reached at SPD before the $i$-th bin, or (B) $k$ number of photons have reached before the $i$-th bin, but none of them has been detected. Therefore, the probability of detecting the first photon in the $i$-th bin, when $n$ photons are present within $T$ time, is given by
@@ IV.4 Time-bin Statistics for Poissonian Photon with Erroneous Devices
Since in the probability distribution of Equation (14), $\frac{T}{N}$ is the length of the time-bins, Equation (17) is identical with Equation (14). Therefore, the probability mass function corresponding to the probability Equation (17) is also given by Equation (15) with $T=Nt_{i}$.
Therefore, these two types of QRNGs, perform similarly in practical scenarios. However, if we can ensure that the source is an actual single-photon source, then the perfect random number can be generated using an external time reference with probability distribution $P_{d}(i)=\frac{d}{N}$ (Ref. Equation (10)). Since the distributions given by Equation (15) are evaluated from the quantum properties of photons, this can be used to test the quantumness in a QRNG.
@@
$f(1)-f(N)<\epsilon$
$\implies\frac{e^{\frac{\mu Td}{N}}-1}{e^{\mu Td}-1}\left(e^{\left(1-\frac{1}{N}\right)\mu Td}-1\right)<\epsilon.$ (18)
We have performed the NIST test Bassham et al. (2010) on the data generated by simulating the probability distribution given by Equation 15. Test results are provided in Appendix B.
V Conclusion
In this article, we assume that photon detection time is available as a sample. We discuss how to determine when the source can be considered as quantum. We suggest two-fold statistical tests. The first one may be an interval estimation or hypothesis test to decide between sub-Poissonian and super-Poissonian. On the other hand, the last one, which is a $\chi^{2}$-test ensures Poisson distribution. This test can be performed for any quantum-photonic device to check the quantumness in it. Then we consider photon statistics for time-bin-based QRNG in practical scenarios. We consider two photon arrival time-based QRNG models, one using external time reference and another without such external reference. We show that, although for ideal devices these two models follow different distributions (one of which is perfectly random), these two QRNGs follow a similar distribution when the photon source is not an actual single photon source. We derive a relation when the photon statistics becomes $\epsilon$-random. We found that by making the product of the expected number of photons ($\mu T$) within an external reference cycle ($T$) and the detector efficiency ($d$) small enough, the randomness of the generated raw random number can be increased. Therefore it would require less post-processing, and it would retain maximum randomness from the quantum source.
In practical cases, the dark count of a detector also contributes to the generation of random numbers. In our discussion, we did not consider the dark count. This work can be extended to see the effect of dark counts on the photon statistics. One may note that our discussions are not meant for black-box testing for commercially available QRNGs. Rather, the physicists and engineers may use these in the QRNG design phase before post-processing, for evaluation, fine-tuning and performance optimization.
VI Acknowledgement
S. Das would like to acknowledge the support received by the Quantum Delta NL KAT-2 project.
Appendix A Relationship between Non-classicality of Light, antibunching of Photons and sub-Poissonian Distributions
The second-order correlation function Fox (2006) of light is defined by
$g^{2}(\tau):=\frac{\langle I(t)I(t+\tau)\rangle}{\langle I(t)\rangle\langle I(t+\tau)\rangle},$ (19)
[…]
$g^{2}(0)$ $=\frac{\langle I(t)^{2}\rangle}{\langle I(t)\rangle^{2}}\geq 1,$ (20)
and $\text{and }g^{2}(\tau)$ $=1\text{ for }\tau\gg\tau_{c},$
$\tau_{c}$ being the coherence time of the source.
Now, consider that a photon stream is sent through a beam splitter and two detectors $D_{0}$ and $D_{1}$ are kept in two output paths. Than the correlation function (19) can be written as
$g^{2}(\tau):=\frac{\langle n_{1}(t)n_{2}(t+\tau)\rangle}{\langle n_{1}(t)\rangle\langle n_{2}(t+\tau)\rangle},$ (21)
where $n_{i}(t)$ is the number of counts registered on detector $i$ at time $t$. If the incoming light consists of a photon stream with long intervals between two consecutive photons, the product $n_{1}(t)n_{2}(t)$ is always 00, giving $g^{2}(0)=0<1$, which is not possible with classical lights. Thus, depending on $g^{2}(t)$, light is distinguished in the following three classes Fox (2006):
  1. (i)
    bunched light, when $g^{2}(0)>1$,
  2. (ii)
    coherent light, when $g^{2}(0)=1$,
  3. (iii)
    antibunched light, when $g^{2}(0)<1$.
Out of these three, antibunched light is undoubtedly non-classical.
The relation between photon count distribution and $g^{2}(\tau)$ is given by Zou and Mandel (1990)
[…]
In the above equation, if $g^{2}(\tau)\leq 1$ for all $\tau$, it implies $\langle(\Delta n)^{2}\rangle<\langle N\rangle$, i.e., a sub-Poissonian statistics, indicating non-classicality of light. However, if $g^{2}(\tau)\leq 1$ for some $\tau$, there is no direct relation between the sub-Poissonian photon count statistic and the antibunching of photons. Xou and Mandel Zou and Mandel (1990) showed that if a source emits photons in two different nonvacuum modes, both having occupation number $n/2$, but different frequencies $\omega_{1}$ and $\omega_{2}$, the second-order correlation function takes the form
$g^{2}(\tau)=\frac{1}{2}\cos(\omega_{1}-\omega_{2})\tau-\frac{1}{n}+1,$ (23)
for $\tau<T$. Also, if $P(N,t,t+T)$ be the distribution of photon count in the time interval from $t$ to $t+T$,
$\langle(\Delta N)^{2}\rangle-\langle N\rangle=\langle N\rangle^{2}\left[\left(\frac{\sin(\omega_{1}-\omega_{2})T/2}{(\omega_{1}-\omega_{2})T/2}\right)^{2}-\frac{1}{n}\right],$ (24)
where $\langle N\rangle$ and $\langle(\Delta N)^{2}\rangle$ are respectively the mean and the variance of the above distribution. Choosing $T=\frac{2\pi}{\omega_{1}-\omega_{2}}$, one can see that $\langle(\Delta N)^{2}\rangle<\langle N\rangle$, i.e., the light is sub-Poissonian although $g^{2}(0)=\frac{3}{2}-\frac{1}{n}>1$ for $n>2$ denoting the light is bunched. This example also shows that although both the antibunching of photons and the sub-Poissonian distribution of photon count indicate the non-classicality of light, they are not identical.
Appendix B Statistical Test of Randomness
As mentioned in Subsection IV.1, the randomness of the random numbers is tested using statistical tests. Different statistical testing suites like NIST Bassham et al. (2010), Dieharder Brown et al. (2014), ENT Walker (2008) etc. are available to serve this purpose. These suites consider different aspects, such as entropy, serial correlation, frequency of values, repetition of similar patterns etc. to perform the test. Some tests also perform Monte Carlo simulation Metropolis and Ulam (1949) to validate the randomness.
The NIST Randomness Testing Suite is an extensive set of statistical tests developed by the National Institute of Standards and Technology (NIST) to evaluate the quality of randomness in random numbers. It is one of the most widely used suites for testing randomness. It contains $15$ different tests to examine different aspects of randomness. Most of the tests have the standard normal or $\chi^{2}$ as reference distribution for their test statistic. Here null hypothesis is taken as the number to be tested is random. For each test, the significance level is considered as $0.01$. If the value of the test statistic for some tests falls in the critical region, or equivalently, calculated $p$-value is less than $0.01$, then the corresponding tests fail rejecting the null hypothesis. A perfectly random number should pass all of the $15$ tests. As mentioned in Section III, the central limit theorem is applicable only for large $n$, tests that use the central limit theorem to determine the reference distribution, have a requirement for the bit stream, that is, the length of the testing sequence, to be large enough (at least $10^{6}$).
Figure 7: NIST result. Bar charts are plotted for (a) $\mu Td=0.05$, (b) $\mu Td=0.10$, (c) $\mu Td=0.15$, (d) $\mu Td=0.20$, (e) $\mu Td=0.25$, (f) $\mu Td=0.30$, (g) $\mu Td=0.35$, (h) $\mu Td=0.40$. For each $\mu Td$, three $1$ GB files are generated by simulating probability distribution (15), and the NIST test has been performed on those files. Different textures denote results corresponding to different files. A test is considered as passed if the corresponding $p$-value is at least $0.01$. Tests that fail for all of the three files are not included in the bar chart. The vertical dashed line denotes $p$-value $=0.01$
@@ Appendix B Statistical Test of Randomness
$0.05$ $10$-$12/15$
$0.10$ $8$-$10/15$
$0.15$ $8/15$
$0.20$ $6$-$7/15$
$0.25$ $6$-$7/15$
$0.30$ $5$-$6/15$
$0.35$ $6/15$
$0.40$ $4$-$5/15$
Table 1: NIST test result for the data simulated with the distribution given by Equation (15). Three files, each containing $1$ GB random number have been used for the test for every $\mu Td$. A test fails if the corresponding $p$-value is found less than $0.01$. The range in the second column arises due to different results from three different files. The number of passed tests decreases with the increment of $\mu Td$. This indicates that as $\mu Td$ increases, the randomness in the generated random number decreases.
We have generated random numbers simulating the probability distribution given by Equation (15) for different $\mu Td$. Three files of size $1$ GB have been generated for every $\mu Td$. Then we perform the NIST tests over the generated numbers. $10$ bit streams of length $10,000,000$ have been considered for the tests. Bar charts of $p$-values from different NIST tests are plotted in Figure 7. The tests for which the $p$-values are found as 00 are not included in the plot. From this plot, one can see that, as $\mu Td$ increases, the number of failed tests also increases indicating a decrement of the randomness in the random number. In Table 1, We tabulate the number of tests our simulated data passed. This table indicates that lower $\mu Td$ values give rise to better random numbers requiring less post-processing.